There are fi ve old ATM machines at Smart Bank. All probability that an ATM breaks down in the course of a day is 0.2. If they are all in working order at the beginning a day, and breakdowns are independent of each other, fi nd the probability that during the day; i. there are no breakdowns ii. four machines break down iii. all machines break down iv. at least three machines break down
1
Expert's answer
2020-04-06T16:35:04-0400

i. This is a binomial distribution with p=0.2, n=5.

"P(X=0)=C^0_5p^0(1-p)^5=(1-0.2)^5=0.32768."


ii. "P(X=4)=C^4_5p^4(1-p)^1=5*0.2^4*(1-0.2)^1=0.0064."


iii. "P(X=5)=C^5_5p^5(1-p)^0=0.2^5=0.00032."


iv. "P(X\ge3)=P(X=3)+P(X=4)+P(X=5)="

"=C^3_5p^3(1-p)^2+C_5^4p^4(1-p)^1+C_5^5p^5(1-p)^0="

"=10*0.2^3*(1-0.2)^2+5*0.2^4+(1-0.2)^1+0.2^5=0.05792."